### v1

# Pranav Gilda    Prakhar Dungarwal    Ariel Thongkham    Elena T. Ajayi    Shaifali Choudhary    Teresa Mondria Terol    Christine Lam    Jessica Pardim Araujo    Megan McFadyen-Mungall    Larry S. Liebovitch    Peter T. Coleman    Harry West    Kate Sieck    Scott Carter

Abstract
We used machine learning and artificial intelligence: 1) to measure levels of peace in countries from news and social media and 2) to develop on-line tools that promote peace by helping users better understand their own media diet.
For news media, we used neural networks to measure levels of peace from text embeddings of on-line news sources. The model, trained on one news media dataset also showed high accuracy when used to analyze a different news dataset.
For social media, such as YouTube, we developed other models to measure levels of social dimensions important in peace using word level (GoEmotions) and context level (Large Language Model) methods.
To promote peace, we note that 71% of people 20-40 years old daily view most of their news through short videos on social media. Content creators of these videos are biased towards creating videos with emotional activation, making you angry to engage you, to increase clicks. We developed and tested a Chrome extension, MirrorMirror, which provides real-time feedback to YouTube viewers about the peacefulness of the media they are watching. Our long term goal is for MirrorMirror to evolve into an open-source tool for content creators, journalists, researchers, platforms, and individual users to better understand the tone of their media creation and consumption and its effects on viewers. Moving beyond simple engagement metrics, we hope to encourage more respectful, nuanced, and informative communication.
Keywords: peace; NLP; machine learning; AI; LLMs.
I Introduction
The content and tone of the language that people use plays an essential role in starting or ending conflicts. “Hate Speech” leads to conflict and violence. Recently, our group and others have been seeking to identify the characteristics of “Peace Speech” that reinforce peaceful behavior [1]. Understanding the importance of the role of language leads to: 1) new ways to measure the levels of peace in countries, regions, or cities; 2) illuminate the social processes that underlie those language differences; and 3) suggest interventions that can promote peace. We show here how modern methods from machine learning and artificial intelligence (AI) can measure levels of peace and design interventions to increase it.
Over the last 25 years our interdisciplinary team and collaborators have used multiple social science methods to study intractable conflicts and sustainable peace [2]. More recently we have used mathematical modeling and data science methods to study and measure peace including:
  • transforming knowledge graphs (also known as causal loop diagrams and concept maps) into sets of ordinary differential equations analyzed as dynamical systems to identify attractors of peace and conflict [3]
  • supervised learning using logistic regression, random forest, support vector machines, decision trees, and embeddings to find the words of highest feature importance that classify lower vs. higher peace countries [1] [4] [5]
  • those words of highest feature importance were clustered into topics using k-means, principal component analysis, large language models, and feedback from workshops with journalists, anthropologists, social psychologists, and poets [6]
  • large language models with retrieval augmented prompt generation (RAG) were used to measure levels of the social processes, positive and negative intergroup reciprocity, in lower and higher peace countries [7]
In this paper, we now report how we extended that previous work to analyze quantitative levels of peace and social dimensions important in peace from on-line news sources and social media such as text transcripts of YouTube videos, using more sophisticated machine learning and AI methods:
  • neural networks using text embeddings
  • Google’s GoEmotions (with the RoBERTa transformer) [8] to measure the emotional tone at the word level
  • large language models (LLMs) such as ChatGPT-3, GPT-4mini [9] and latest others from Google Gemini to measure meanings at the context level
We also report that we developed a real-time tool to foster peace. “We become what we behold” [10]. 71% of people ages 16 to 40 get their news daily from social media, especially Facebook and YouTube [11]. Content creators on those sites create videos to increase emotional activation, especially anger, to increase engagement and clicks. We sought to apply our research findings to see how we can go beyond simple engagement metrics to encourage more respectful, nuanced, and informative communication.
We developed a Chrome extension, MirrorMirror, that provides real-time feedback to YouTube viewers on the peacefulness of the videos they are watching. This was done through Human Centered Design (HCD) using iterative feedback from our team members, students in a design class, and a broader set of people to design the User Experience / User Interface (UX/UI). Our long term goal is for MirrorMirror to evolve into an open-source tool for content creators, journalists, researchers, platforms, and individual users to better understand the tone of their media creation and consumption.
We will make our data/code available upon acceptance on our GitHub repository.
II Methods: News Data
II-A Strategy
There are many surveys that measure levels of peace in a country conducted by international organizations, national governmental agencies, non-profit groups, and for-profit consulting companies [1]. Thus, we can use that data to tag samples of text from different countries with their level of peace to form training and testing sets for supervised learning. We also tested the accuracy of models trained on one dataset analyzing another dataset of different countries.
II-B Neural Network Classification of Peace
II-B1 Data Preparation
A large-scale training dataset was constructed from the News on the Web (NOW) corpus, comprising approximately 700,000 news articles from 18 countries. Each article included metadata specifying its country of origin and the associated peace level. High- and low-peace labels were assigned following the method described by Liebovitch et al. [1], which maps linguistic features of national media to established peace indices.
To prepare the data, n-gram preprocessing was applied to capture multi-word contextual patterns and reduce sparsity. The processed text was then embedded using the OpenAI text-embedding-3-small model, which converts each article into a 1,536-dimensional vector representation encoding its semantic meaning. Of the 18 countries represented, 10 were identified as high-peace, resulting in a balanced distribution for unbiased model training and evaluation.
II-B2 Model Architectures
Three neural architectures were developed to perform binary classification of articles as high-peace or low-peace:
  • CNN Model: Two convolutional layers (64 and 32 filters; kernel size = 3, ReLU activation) were followed by a flattening layer and two dense layers (128 and 64 units) with a 0.3 dropout rate for regularization and a sigmoid output. This configuration was designed to extract localized semantic patterns within the embeddings, capturing phrases or n-grams indicative of peaceful discourse.
  • Feed-Forward Model: A fully connected network of four dense layers (512, 256, 128, and 64 units), each followed by 0.3 dropout, concluded with a sigmoid output layer. This deeper architecture enabled the model to learn global semantic relationships across articles, representing broader conceptual differences in peace-related language.
  • Revised CNN Model: The revised design incorporated a max-pooling layer (pool size = 2) between convolutional layers, enabling dimensionality reduction and improved generalization. It retained the two fully connected layers (128 and 64 units) from the original CNN, followed by a sigmoid classifier.
These architectures were selected to compare the performance of spatially sensitive versus fully connected learning strategies on high-dimensional text embeddings.
II-B3 Training Procedure
The dataset was divided into 80% training and 20% testing subsets using a fixed random seed to ensure reproducibility. For the convolutional models, each input sample was represented as a sequence of 1,536 elements with a single feature channel, while the feed-forward network received the same 1,536-dimensional vectors in flattened form. All models were trained for 10 epochs with a batch size of 32, using the Adam optimizer (learning rate = 0.001) and binary cross-entropy loss. Accuracy served as the primary evaluation metric.
Training histories were recorded to monitor loss and accuracy over epochs, assessing both convergence and generalization on the test set. Dropout regularization (0.3) was applied to mitigate overfitting. No separate validation set was used, as the objective was not hyperparameter tuning but rather establishing a consistent baseline for the networks’ ability to classify high-peace versus low-peace texts. All models were trained on GPU-enabled hardware using TensorFlow 2.15 and Python 3.10.
Data Collection NOW corpus, 700K articlesPeace Label Assignment High/low-peace countries [1]N-gram PreprocessingText Embedding OpenAI, 1,536-dim vectorsTrain/Test Split (80/20)CNNFeed-ForwardRevised CNNModel Training Adam, 10 epochs, batch 32Evaluation Accuracy & Loss
Figure 1: Neural network training and evaluation pipeline for peace classification.
III Results: News Data
III-A Evaluation of Neural Networks for Peace Classification
III-A1 Model Performance on Test Data
Across the three network architectures, all models achieved high performance on the held-out test set from the NOW dataset. The convolutional neural network (CNN) achieved a test accuracy of 97.24%, the fully connected feed-forward network reached 97.48%, and the revised CNN achieved 96.99%. These results indicate that the models effectively learned to distinguish linguistic patterns associated with high- and low-peace discourse within structured news text.
Although all three models performed comparably, the feed-forward network’s slightly higher accuracy aligns with its architectural advantage: its dense layers integrate global semantic relationships across all embedding dimensions. This capability is particularly beneficial in a binary classification task such as peace prediction, where subtle distinctions often depend on distributed semantic cues across the entire embedding space. By contrast, the CNNs—while effective at detecting local n-gram-like patterns—may overlook long-range dependencies essential for capturing the overall semantic and emotional tone of written language. Nonetheless, the CNNs’ strong performance demonstrates that both local and global text features contributed meaningfully to the classification of peace-related language.
III-A2 Cross-Dataset Generalization
To assess the generalizability of the trained models, they were tested on an independent corpus—the Capstone Peace Speech dataset comprising 600,000 news articles from 16 countries (eight high-peace and eight low-peace). Each article was labeled according to established peace indices such as the Global Peace Index and the Positive Peace Index. Despite differences in collection methodology and source distribution, the models show substantial cross-dataset transfer in the per-article predictions: 72.81% for the CNN, 72.47% for the feed-forward network, and 69.91% for the revised CNN. When grouped and averaged at the country level, all three networks successfully classified every country in accordance with its peace label. This result suggests that the linguistic features learned from the NOW dataset capture generalizable distinctions in structure, tone, and framing that align with broader indicators of national peacefulness.
TABLE I: Neural Network Performance Across Datasets
Model NOW (%) Capstone (%)
CNN 97.24 72.81
Feed-Forward 97.48 72.47
Revised CNN 96.99 69.91
III-A3 Comparison Between Network Architectures
The per-article Capstone evaluation shows that the original CNN transfers slightly better than the revised CNN and performs comparably to the feed-forward network. A likely reason is that the original convolutional design, without the added pooling layer, retains fine-grained local patterns—short semantic or tonal cues that may remain stable across domains. The feed-forward network, meanwhile, leverages global relationships across all embedding dimensions. Overall, the strong performance across models indicates that peace-related linguistic markers exist at multiple levels of representation and can be effectively captured by both convolutional and fully connected architectures.
III-A4 Cross-Domain Generalization to YouTube Transcripts
To explore whether the trained models could generalize beyond formal written text, they were applied to a smaller dataset of 22 YouTube transcripts drawn from five major news organizations: The New York Times, CNN, NBC, The Washington Post, and Breitbart News. However, when evaluated on this dataset, all networks exhibited severe overgeneralization, classifying 95–100% of the videos as “high-peace.”
This outcome suggests a transfer failure from written to spoken media. The linguistic characteristics of video journalism differ substantially from print media—not only in syntax and register but also in the use of immediacy, narrative framing, and emotional activation. Video transcripts often include conversational fillers, hesitations, and speaker attributions that are absent from formal writing. These stylistic and pragmatic differences likely disrupted the networks’ ability to apply the same semantic cues used to identify peace-related language in articles. This finding motivated our subsequent use of emotion-based and large language model (LLM) techniques to better capture nuanced indicators of peace in multimodal and social media contexts.
IV Methods: YouTube Data
IV-A Strategy
The news data models were trained by tagging news data with the levels of peace from external traditional peace indices. There is no similar external peace measures to tag YouTube videos. Therefore, we based our YouTube models both on 1) feedback that we received from peace experts and journalists reviewing our previous machine learning results and 2) on social dimensions identified by 59 studies over the last 40 years as important measures of the level of peace, including studies on: integrative complexity, political framing meta-analysis, tightness/looseness, and many others [12][13][14][15]. These dimensions are:
compassion — contempt
news — opinion
prevention — promotion
order — creativity
nuance — simplistic
IV-B Analytical Models for YouTube Transcripts
The failure of the news-trained networks demonstrated that peace in video content is not defined by specific keywords (topics) but by the manner in which those topics are discussed (tone and framing). While Neural Networks excelled at detecting topic-based patterns in news, they failed to capture the conversational volatility of video. We therefore pivoted to models designed to detect emotional tone (GoEmotions) and contextual framing (LLMs).
IV-C Google GoEmotions
Our first approach utilized Google’s GoEmotions, a RoBERTa-based transformer model trained on a large corpus of Reddit comments to classify text into 28 emotion categories [8]. We hypothesized that we could map these fine-grained emotions to our five broader peace dimensions. However, this approach revealed several significant limitations:
  • Limited Context Window: The model analyzes text in small, discrete chunks (e.g., sentence-by-sentence), lacking a broader, holistic understanding of the video’s full narrative arc. While we aggregated scores at the paragraph level, this post-processing step could not fix the underlying lack of context.
  • High Neutrality Baseline: The model, trained on a different domain (Reddit comments), frequently classified large portions (40-70%) of the YouTube transcripts as “neutral.” This high baseline score obscured the more subtle emotional cues indicative of bias or nuance, limiting the model’s utility.
  • The Averaging-Out Problem: A critical failure mode of simple aggregation is its inability to capture emotional volatility. For example, a transcript that expressed strong contempt for one topic (e.g., immigrants) and later expressed strong praise for another (e.g., a policy) would have its scores averaged, misleading the results. This process effectively erased the very nuance we intended to measure.
  • Deterministic Weighted Mapping: We attempted a heuristic approach by mapping GoEmotions outputs to a scalar valence score (e.g., joy/admiration = +1.0, anger/disgust = -1.0, neutral = 0.0). However, this deterministic mapping failed to correlate with the dimensions of the peace because it ignored the context.
These limitations made it clear that while the GoEmotions scores were valuable as a quantitative feature, they were insufficient as a standalone analyzer for our complex peace dimensions.
Motivation for a Synthesis Model: This led us to hypothesize that a more robust solution would be a multi-stage process. We needed a model that could not only see the quantitative emotional data from GoEmotions but also interpret it within the full context of the entire transcript.
This motivated our pivot to Large Language Models (LLMs) as a synthesis tool. We theorized that LLMs, while potentially subjective when used in isolation, could be ”grounded” by providing the GoEmotions scores as a structured, quantitative input. This dual-input approach, giving the LLM both the text and its emotional profile—was designed to improve trustworthiness, observability, and provide a tighter feedback loop for our team to evaluate the model’s reasoning.
IV-D Large Language Models (LLMs)
In parallel with the GoEmotions analysis, we explored the capabilities of Large Language Models (LLMs) to score transcripts along the same peace dimensions. Our methodology evolved through several stages of prompt engineering, moving from a general-purpose prompt to a final, sophisticated model that synthesized textual data with quantitative emotional analysis.
Initial Prompting (News Media Framework)
We initially adapted a comprehensive prompt designed for a related project analyzing formal news articles (PIR/NIR concepts). This prompt, tested on models such as OpenAI’s GPT-3 and GPT-4-mini, was highly detailed. It provided a role for the AI (“media analyst”), a +5 to -5 scoring scale, and multiple, detailed examples for each of the five dimensions.
While effective for structured news text, this prompt proved overly complex and ill-suited for the conversational, faster-paced, and stylistically different nature of YouTube video transcripts. The extensive rules and examples did not generalize well to the broader domains, leading to inconsistent and unreliable scoring.
Iterative Refinement (Simplistic-Generalizable Prompt)
Recognizing the limitations of the initial prompt, we engaged in an iterative engineering process to develop a version optimized for YouTube transcripts. The primary goals were to simplify the instructions leading to better generalizability, maintain the core 5-dimensional analysis.
IV-E Human Feedback
To establish a rigorous benchmark for evaluating our automated approaches, we collaborated with human experts specializing in peace research and conflict studies. Our team of peace researchers, including social psychologists, anthropologists, and conflict resolution specialists from Columbia University’s Morton Deutsch International Center for Cooperation and Conflict Resolution, independently evaluated a gold standard set of 52 YouTube videos.
Each expert scored videos across all five dimensions using the same 1-5 scale employed by our computational models. Videos were selected to represent a diverse range of content types, including political commentary, news analysis, social issues discussion, and cultural critique.
Inter-rater reliability analysis validated the robustness of this gold standard, yielding pairwise correlations exceeding $r=0.93$ for three dimensions (News, Compassion, Order) and 100% agreement within a single-point margin. This high consensus confirms that our peace dimensions are observable and measurable constructs, distinct from random subjective noise. We aggregated human scores by averaging across all available raters for each video-dimension combination, with sample sizes ranging from N=32 to N=47 depending on missing data patterns. The resulting gold standard dataset provides mean scores, standard deviations, and distributional statistics that serve as ground truth for evaluating model performance.
TABLE II: Human Expert Annotation Statistics
Dimension N Mean SD Min Max Median
Nuance-Simplistic 37 2.76 1.06 1.0 4.83 3.0
Order-Creativity 32 2.81 1.38 1.0 5.0 3.0
Prevention-Promotion 35 2.64 1.32 1.0 5.0 2.5
Compassion-Contempt 47 2.89 0.94 1.0 5.0 3.0
Opinion-News 34 3.06 1.43 1.0 5.0 3.4
V Results: YouTube Data
The AI models that we used are shown in Fig. 2, we compared the predicted values from those models analysis of the transcripts of YouTube videos with those reported by our human coders on five social dimensions.
Summary of YouTube Results As shown in Fig. 2, the values of compassion—contempt predicted by the models, such as Google’s GoEmotions, based on the valence of individual words, even if enhanced by the transformer RoBERTa, were only weakly correlated, $r\approx 0.18$, with those of our human coders. As we gave that social dimension the highest priority in our first designs of MirrorMirror, we then explored the use of LLMs.
As shown in Fig. 3, modern large language models can reliably evaluate multidimensional aspects of peace journalism in online video content, achieving correlations (up to $r=0.773$) comparable to human inter-rater agreement. This suggests that AI systems can meaningfully scale and augment human media analysis. Gemini 3 Pro Preview achieved the highest correlations across all five dimensions, marking a critical +0.317+0.317 improvement over previous iterations in the challenging Nuance dimension. This suggests that the latest reasoning models are beginning to bridge the gap in understanding context-heavy ”gray areas” where earlier models struggled.
Figure 2: Pearson correlation coefficient $r$ between the human coders and predictions of the AI models on the dimension: compassion—contempt. We used the models: Gemini 3 Pro Preview, Gemini 2.5 Flash, GPT-5.1, GPT-4o, RoBERTa, and GoEmotions. +R indicates hybrid models with RoBERTa.
Figure 3: Pearson correlation coefficient $r$ between the human coders and predictions of the AI models for all 5 social science dimensions.
LLMs Understand Context Beyond Surface Emotion Signals. Across all dimensions, LLMs substantially outperformed simpler emotion-only baselines, indicating that peace assessment requires contextual reasoning—tone, framing, rhetorical structure, and implicit messaging—rather than raw sentiment. The relatively weaker performance on the Nuance dimension reflects a fundamental challenge: recognizing multiple perspectives may require deeper external knowledge, pointing toward the value of retrieval-augmented or fact-aware models.
Model Architecture and Peace-Metric Alignment. This consistent performance advantage suggests that the Gemini architecture may be better aligned with the nuances of sociological scoring tasks, underscoring the importance of empirical evaluation over vendor expectations on domain-specific tasks rather than relying on general benchmarks. Emotion-augmented prompting yielded mixed effects, suggesting that additional signals do not automatically improve reasoning.
Multimodality and Agents Limitations such as dataset size, lack of multimodal cues, and absence of causal evidence highlight clear next steps: integrating audio-visual features, exploring multi-agent assessment pipelines, developing smaller fine-tuned models for edge deployment, and extending across platforms such as TikTok and Twitter.
VI Discussion
The next steps in developing our MirrorMirror chrome extension are: 1) to use more extensive human feedback to compare the accuracy of different computational methods to measure the social dimensions in the YouTube videos, 2) to expand our testing of the UX/UI interface across a broader group of people, 3) to provide users information about how the videos they watch change over time, and 4) to do behavioral testing to measure the effects of MirrorMirror on users choice of videos and if those choices lead to improved behaviors with others.
VII Conclusions
We have shown
  • how levels of peace can be measured from news text sources using data-driven supervised machine learning
  • how dimensions of social factors important in peaceful cultures, as identified by human experts and hypothesis-driven social science studies, can be measured from social media text using word and context level AI tools
We have used those results
  • to develop a chrome extension that can provide users useful feedback to increase their awareness of how they are influenced by their own media diet. Our initial user interface is shown in Fig. 4. Perhaps, such self-awareness may be a helpful tool in turning down the destructive, polarizing, heat in social media communications.
Figure 4: Prototype of the MirrorMirror user interface.

### v2

@@
-Pranav Gilda    Prakhar Dungarwal    Ariel Thongkham    Elena T.
-Ajayi    Shaifali Choudhary    Teresa Mondria Terol    Christine Lam    Jessica Pardim Araujo    Megan McFadyen-Mungall    Larry S.
-Liebovitch    Peter T.
-Coleman    Harry West    Kate Sieck    Scott Carter
+
 Abstract
-We used machine learning and artificial intelligence: 1) to measure levels of peace in countries from news and social media and 2) to develop on-line tools that promote peace by helping users better understand their own media diet.
-For news media, we used neural networks to measure levels of peace from text embeddings of on-line news sources.
+We used machine learning and artificial intelligence: 1) to measure levels of peace in countries from news and social media and 2) to develop online tools that promote peace by helping users better understand their own media diet.
+For news media, we used neural networks to measure levels of peace from text embeddings of online news sources.
 The model, trained on one news media dataset also showed high accuracy when used to analyze a different news dataset.
 For social media, such as YouTube, we developed other models to measure levels of social dimensions important in peace using word level (GoEmotions) and context level (Large Language Model) methods.
-To promote peace, we note that 71% of people 20-40 years old daily view most of their news through short videos on social media.
+To promote peace, we note that 71% of people 16-40 years old daily view most of their news through short videos on social media.
 Content creators of these videos are biased towards creating videos with emotional activation, making you angry to engage you, to increase clicks.
@@
   • large language models with retrieval augmented prompt generation (RAG) were used to measure levels of the social processes, positive and negative intergroup reciprocity, in lower and higher peace countries [7]
-In this paper, we now report how we extended that previous work to analyze quantitative levels of peace and social dimensions important in peace from on-line news sources and social media such as text transcripts of YouTube videos, using more sophisticated machine learning and AI methods:
+In this paper, we now report how we extended that previous work to analyze quantitative levels of peace and social dimensions important in peace from online news sources and social media such as text transcripts of YouTube videos, using more sophisticated machine learning and AI methods:
   • neural networks using text embeddings
@@
 Thus, we can use that data to tag samples of text from different countries with their level of peace to form training and testing sets for supervised learning.
-We also tested the accuracy of models trained on one dataset analyzing another dataset of different countries.
+We also tested the accuracy of models trained on one dataset analyzing another dataset of different countries, see Fig.
+1.
 II-B Neural Network Classification of Peace
@@
 Figure 1: Neural network training and evaluation pipeline for peace classification.
-III Results: News Data
-III-A Evaluation of Neural Networks for Peace Classification
-III-A1 Model Performance on Test Data
-Across the three network architectures, all models achieved high performance on the held-out test set from the NOW dataset.
-The convolutional neural network (CNN) achieved a test accuracy of 97.24%, the fully connected feed-forward network reached 97.48%, and the revised CNN achieved 96.99%.
-These results indicate that the models effectively learned to distinguish linguistic patterns associated with high- and low-peace discourse within structured news text.
-Although all three models performed comparably, the feed-forward network’s slightly higher accuracy aligns with its architectural advantage: its dense layers integrate global semantic relationships across all embedding dimensions.
-This capability is particularly beneficial in a binary classification task such as peace prediction, where subtle distinctions often depend on distributed semantic cues across the entire embedding space.
-By contrast, the CNNs—while effective at detecting local n-gram-like patterns—may overlook long-range dependencies essential for capturing the overall semantic and emotional tone of written language.
-Nonetheless, the CNNs’ strong performance demonstrates that both local and global text features contributed meaningfully to the classification of peace-related language.
-III-A2 Cross-Dataset Generalization
-To assess the generalizability of the trained models, they were tested on an independent corpus—the Capstone Peace Speech dataset comprising 600,000 news articles from 16 countries (eight high-peace and eight low-peace).
-Each article was labeled according to established peace indices such as the Global Peace Index and the Positive Peace Index.
-Despite differences in collection methodology and source distribution, the models show substantial cross-dataset transfer in the per-article predictions: 72.81% for the CNN, 72.47% for the feed-forward network, and 69.91% for the revised CNN.
-When grouped and averaged at the country level, all three networks successfully classified every country in accordance with its peace label.
-This result suggests that the linguistic features learned from the NOW dataset capture generalizable distinctions in structure, tone, and framing that align with broader indicators of national peacefulness.
-TABLE I: Neural Network Performance Across Datasets
-Model NOW (%) Capstone (%)
-CNN 97.24 72.81
-Feed-Forward 97.48 72.47
-Revised CNN 96.99 69.91
-III-A3 Comparison Between Network Architectures
-The per-article Capstone evaluation shows that the original CNN transfers slightly better than the revised CNN and performs comparably to the feed-forward network.
-A likely reason is that the original convolutional design, without the added pooling layer, retains fine-grained local patterns—short semantic or tonal cues that may remain stable across domains.
-The feed-forward network, meanwhile, leverages global relationships across all embedding dimensions.
-Overall, the strong performance across models indicates that peace-related linguistic markers exist at multiple levels of representation and can be effectively captured by both convolutional and fully connected architectures.
-III-A4 Cross-Domain Generalization to YouTube Transcripts
-To explore whether the trained models could generalize beyond formal written text, they were applied to a smaller dataset of 22 YouTube transcripts drawn from five major news organizations: The New York Times, CNN, NBC, The Washington Post, and Breitbart News.
-However, when evaluated on this dataset, all networks exhibited severe overgeneralization, classifying 95–100% of the videos as “high-peace.”
-This outcome suggests a transfer failure from written to spoken media.
-The linguistic characteristics of video journalism differ substantially from print media—not only in syntax and register but also in the use of immediacy, narrative framing, and emotional activation.
-Video transcripts often include conversational fillers, hesitations, and speaker attributions that are absent from formal writing.
-These stylistic and pragmatic differences likely disrupted the networks’ ability to apply the same semantic cues used to identify peace-related language in articles.
-This finding motivated our subsequent use of emotion-based and large language model (LLM) techniques to better capture nuanced indicators of peace in multimodal and social media contexts.
-IV Methods: YouTube Data
-IV-A Strategy
+III Methods: YouTube Data
+III-A Strategy
 The news data models were trained by tagging news data with the levels of peace from external traditional peace indices.
-There is no similar external peace measures to tag YouTube videos.
+There are no similar external peace measures to tag YouTube video transcripts.
 Therefore, we based our YouTube models both on 1) feedback that we received from peace experts and journalists reviewing our previous machine learning results and 2) on social dimensions identified by 59 studies over the last 40 years as important measures of the level of peace, including studies on: integrative complexity, political framing meta-analysis, tightness/looseness, and many others [12][13][14][15].
@@
 nuance — simplistic
-IV-B Analytical Models for YouTube Transcripts
+III-B Analytical Models for YouTube Transcripts
 The failure of the news-trained networks demonstrated that peace in video content is not defined by specific keywords (topics) but by the manner in which those topics are discussed (tone and framing).
@@
 We therefore pivoted to models designed to detect emotional tone (GoEmotions) and contextual framing (LLMs).
-IV-C Google GoEmotions
+III-C Google GoEmotions
 Our first approach utilized Google’s GoEmotions, a RoBERTa-based transformer model trained on a large corpus of Reddit comments to classify text into 28 emotion categories [8].
@@
 This dual-input approach, giving the LLM both the text and its emotional profile—was designed to improve trustworthiness, observability, and provide a tighter feedback loop for our team to evaluate the model’s reasoning.
-IV-D Large Language Models (LLMs)
+III-D Large Language Models (LLMs)
 In parallel with the GoEmotions analysis, we explored the capabilities of Large Language Models (LLMs) to score transcripts along the same peace dimensions.
@@
 The primary goals were to simplify the instructions leading to better generalizability, maintain the core 5-dimensional analysis.
-IV-E Human Feedback
+III-E Human Feedback
 To establish a rigorous benchmark for evaluating our automated approaches, we collaborated with human experts specializing in peace research and conflict studies.
@@
 We aggregated human scores by averaging across all available raters for each video-dimension combination, with sample sizes ranging from N=32 to N=47 depending on missing data patterns.
-The resulting gold standard dataset provides mean scores, standard deviations, and distributional statistics that serve as ground truth for evaluating model performance.
-TABLE II: Human Expert Annotation Statistics
+The resulting gold standard dataset provides mean scores, standard deviations, and distributional statistics that serve as ground truth for evaluating model performance, as shown in Table I.
+TABLE I: Human Expert Annotation Statistics
 Dimension N Mean SD Min Max Median
@@
 Opinion-News 34 3.06 1.43 1.0 5.0 3.4
+IV Results: News Data
+IV-A Evaluation of Neural Networks for Peace Classification
+IV-A1 Model Performance on Test Data
+Across the three network architectures, all models achieved high performance on the held-out test set from the NOW dataset, as shown in Table II, The convolutional neural network (CNN) achieved a test accuracy of 97.24%, the fully connected feed-forward network reached 97.48%, and the revised CNN achieved 96.99%.
+These results indicate that the models effectively learned to distinguish linguistic patterns associated with high- and low-peace discourse within structured news text.
+Although all three models performed comparably, the feed-forward network’s slightly higher accuracy aligns with its architectural advantage: its dense layers integrate global semantic relationships across all embedding dimensions.
+This capability is particularly beneficial in a binary classification task such as peace prediction, where subtle distinctions often depend on distributed semantic cues across the entire embedding space.
+By contrast, the CNNs—while effective at detecting local n-gram-like patterns—may overlook long-range dependencies essential for capturing the overall semantic and emotional tone of written language.
+Nonetheless, the CNNs’ strong performance demonstrates that both local and global text features contributed meaningfully to the classification of peace-related language.
+IV-A2 Cross-Dataset Generalization
+To assess the generalizability of the trained models, they were tested on an independent corpus—the Capstone Peace Speech dataset comprising 600,000 news articles from 16 countries (eight high-peace and eight low-peace).
+Each article was labeled according to established peace indices such as the Global Peace Index and the Positive Peace Index.
+Despite differences in collection methodology and source distribution, the models show substantial cross-dataset transfer in the per-article predictions: 72.81% for the CNN, 72.47% for the feed-forward network, and 69.91% for the revised CNN.
+When grouped and averaged at the country level, all three networks successfully classified every country in accordance with its peace label.
+This result suggests that the linguistic features learned from the NOW dataset capture generalizable distinctions in structure, tone, and framing that align with broader indicators of national peacefulness.
+TABLE II: Neural Network Performance Across Datasets
+Model NOW (%) Capstone (%)
+CNN 97.24 72.81
+Feed-Forward 97.48 72.47
+Revised CNN 96.99 69.91
+IV-A3 Comparison Between Network Architectures
+The per-article Capstone evaluation shows that the original CNN transfers slightly better than the revised CNN and performs comparably to the feed-forward network.
+A likely reason is that the original convolutional design, without the added pooling layer, retains fine-grained local patterns—short semantic or tonal cues that may remain stable across domains.
+The feed-forward network, meanwhile, leverages global relationships across all embedding dimensions.
+Overall, the strong performance across models indicates that peace-related linguistic markers exist at multiple levels of representation and can be effectively captured by both convolutional and fully connected architectures.
+IV-A4 Cross-Domain Generalization to YouTube Transcripts
+To explore whether the trained models could generalize beyond formal written text, they were applied to a smaller dataset of 22 YouTube transcripts drawn from five major news organizations: The New York Times, CNN, NBC, The Washington Post, and Breitbart News.
+However, when evaluated on this dataset, all networks exhibited severe overgeneralization, classifying 95–100% of the videos as “high-peace.”
+This outcome suggests a transfer failure from written to spoken media.
+The linguistic characteristics of video journalism differ substantially from print media—not only in syntax and register but also in the use of immediacy, narrative framing, and emotional activation.
+Video transcripts often include conversational fillers, hesitations, and speaker attributions that are absent from formal writing.
+These stylistic and pragmatic differences likely disrupted the networks’ ability to apply the same semantic cues used to identify peace-related language in articles.
+This finding motivated our subsequent use of emotion-based and large language model (LLM) techniques to better capture nuanced indicators of peace in multimodal and social media contexts.
 V Results: YouTube Data